Equation of circle whose radius is 4 unit, centre is on x-axis and passes through origin--Explain????

Grade:11

1 Answers

Vikas TU
7 years ago
Let the centre be (a,0)
then from the distance formulae we get,
root(a^2 +0^2) = 4
or
a^2 = 16
a = +4 or -4
Hence, eqn. of circle,
(x-4)^2 + y^2 = 16